# Direct Sampling from Conditional Distributions by Sequential Maximum Likelihood Estimations

### v2

@@ 4 Discussion
Previous implementations of the direct sampling algorithm [14, 16] have enabled exact sampling, but depending on the model, the computational burden makes it challenging to put into practical use. In this study, we proposed the approximate algorithm applicable to any log-affine model with the expense of the bias, that asymptotically disappears. In contrast to deriving the Markov basis in the Metropolis algorithm or the connection matrix of the holonomic ideal in the exact sampling, we can implement the approximate sampling algorithm straightforwardly since it demands no prior investigations; it only needs existing methods to compute the MLE.
There is room for improvement in the approximate direct sampling algorithm proposed in this paper. In terms of accuracy, it can approach the UMVUE by reducing the bias of the MLE of the expected count, and in terms of the speed of computing the MLE, Newton’s method can be faster than the generalized iterative proportional scaling. However, these are typical tasks in statistical analysis when applying an algorithm to each model and are beyond the scope of this paper, which aims to propose an approximate sampling algorithm in its most general form.
@@ Acknowledgements
This work was supported in part by JSPS KAKENHI Grant Numbers 20K03742 and 24K06876.

### v3

@@ 4 Discussion
Previous implementations of the direct sampling algorithm [15, 17] have enabled exact sampling, but depending on the model, the computational burden makes it challenging to put into practical use. In this study, we proposed the approximate algorithm applicable to any log-affine model with the expense of the bias, that asymptotically disappears. In contrast to deriving the Markov basis in the Metropolis algorithm or the connection matrix of the holonomic ideal in the exact sampling, we can implement the approximate sampling algorithm straightforwardly since it demands no prior investigations; it only needs existing methods to compute the MLE. Furthermore, as shown in Section 3.1 for a non-independence model of two-way contingency tables, the approximate direct sampling algorithm achieves more accurate sampling than the Metropolis algorithm for a fixed number of tables in the same order of time, despite the bias coming from using the MLE to compute the transition probability.
In the worst case, approximate direct sampling for the no-three-way interaction model in Section 3.1 took 36 times the computational time of the Metropolis algorithm. While it is true that approximate direct sampling is computationally expensive, we should note that a Markov basis for the no-three-way interaction model is known and given. If a Markov basis is unknown for a model and cannot be calculated, the Metropolis algorithm cannot be applied. In contrast, the approximate direct sampling is always feasible as long as maximum likelihood estimation is possible.
The approximate direct sampling algorithm can be improved. In terms of accuracy, the MLE approaches the UMVUE by reducing the bias. The UMVUE could be used in the later steps of the algorithm where the total number of counts becomes small. For computing the MLE, Newton’s method is likely faster than the generalized iterative proportional scaling. However, these are typical tasks in statistical analysis when applying an algorithm to each model and are beyond the scope of this paper, which aims to propose an approximate sampling algorithm in its most general form.
@@ Acknowledgements
The author thanks the referees for their valuable comments on improving the presentation. He also thanks a referee for pointing out the mistake in the original version of Theorem 2.9 and suggesting improvements. This work was supported in part by JSPS KAKENHI Grant Numbers 20K03742 and 24K06876.
Appendix
We show that the rational MLE and the UVMUE coincide in the two-way contingency table of the quasi-independence model that appeared in Example 2.12. If we fix $u_{13}$ and $u_{31}$, we have the $2\times 2$ subtable:
$\begin{array}[]{cc|c}u_{11}&u_{12}&u_{1\cdot}-u_{13}\\ u_{21}&u_{22}&u_{2\cdot}-u_{23}\\ \hline\cr u_{\cdot 1}-u_{31}&u_{\cdot 2}-u_{32}&u_{1\cdot}+u_{2\cdot}-u_{\cdot 3}\end{array},$
where $u_{1\cdot}+u_{2\cdot}-u_{\cdot 3}=u_{\cdot 1}+u_{\cdot 2}-u_{3\cdot}$, $u_{3\cdot}=u_{31}+u_{32}$, and $u_{\cdot 3}=u_{13}+u_{23}$. The probability mass function of the hypergeometric distribution gives
$\sum_{u_{11}}\frac{1}{u_{11}!u_{12}!u_{21}!u_{22}!}=\frac{(u_{1\cdot}+u_{2\cdot}-u_{\cdot 3})!}{(u_{1\cdot}-u_{13})!(u_{2\cdot}-u_{23})!(u_{\cdot 1}-u_{31})!(u_{\cdot 2}-u_{32})!}.$
Therefore,
$Z_{A}(b;1)$ $=\sum_{u_{13},u_{31}}\frac{1}{u_{13}!u_{23}!u_{31}!u_{32}!}\frac{(u_{1\cdot}+u_{2\cdot}-u_{\cdot 3})!}{(u_{1\cdot}-u_{13})!(u_{2\cdot}-u_{23})!(u_{\cdot 1}-u_{31})!(u_{\cdot 2}-u_{32})!}$
$=(u_{1\cdot}+u_{2\cdot}-u_{\cdot 3})!\sum_{u_{13}}\frac{1}{u_{13}!(u_{1\cdot}-u_{13})!(u_{\cdot 3}-u_{13})!(u_{2\cdot}-u_{\cdot 3}+u_{13})!}$
$~~\times\sum_{u_{31}}\frac{1}{u_{31}!(u_{\cdot 1}-u_{31})!(u_{3\cdot}-u_{31})!(u_{\cdot 2}-u_{3\cdot}+u_{31})!}$
$=\frac{(u_{1\cdot}+u_{2\cdot}-u_{\cdot 3})!}{u_{1\cdot}!u_{\cdot 3}!(u_{2\cdot}-u_{\cdot 3})!u_{\cdot 1}!u_{3\cdot}!(u_{\cdot 2}-u_{3\cdot})!}$
$~~\times{}_{2}F_{1}(-u_{1\cdot},-u_{\cdot 3},u_{2\cdot}-u_{\cdot 3}+1)\times{}_{2}F_{1}(-u_{\cdot 1},-u_{3\cdot},u_{\cdot 2}-u_{3\cdot}+1)$
$=\frac{(u_{1\cdot}+u_{2\cdot})!(u_{\cdot 1}+u_{\cdot 2})!}{u_{\cdot 1}!u_{\cdot 2}!u_{\cdot 3}!u_{1\cdot}!u_{2\cdot}!u_{3\cdot}!(u_{\cdot 1}+u_{\cdot 2}-u_{3\cdot})!},$
where in the last equality we used the Gauss hypergeometric theorem
$_{2}F_{1}(\alpha,\beta;\gamma;1)=\frac{\Gamma(\gamma)\Gamma(\gamma-\alpha-\beta)}{\Gamma(\gamma-\alpha)\Gamma(\gamma-\beta)},\quad{\rm Re}(\gamma)>{\rm Re}(\alpha+\beta),\quad-\gamma\not\in\mathbb{N}$
for $\alpha,\beta,\gamma\in\mathbb{C}$. Then, the UMVUE is
$\tilde{\mu}_{13}(b;1)=\frac{Z_{A}(b-a_{13};1)}{Z_{A}(b;1)}=\frac{u_{1\cdot}u_{\cdot 3}}{u_{1\cdot}+u_{2\cdot}},$
where $a_{13}$ is the column vector of $A$ corresponding to the first row and the third column. The UMVUE coincides with the rational MLE in Example 2.5 of [4].